# Regularized least squares learning with heavy-tailed noise is minimax optimal

### v2

@@ A.3.1 Bounding the variance $I_{\pi}\widehat{f}_{\alpha}-I_{\pi}f_{\alpha}$
$=\underbrace{I_{\pi}(\widehat{\mathcal{C}}_{\pi}+\alpha\operatorname{Id}_{\mathcal{H}})^{-1}\mathopen{}\mathclose{{}\left(\widehat{\Upsilon}-\widehat{\mathcal{C}}_{\pi}f_{\alpha}}\right)}_{=:I_{1}}+\underbrace{\alpha I_{\pi}(\widehat{\mathcal{C}}_{\pi}+\alpha\operatorname{Id}_{\mathcal{H}})^{-1}f_{\alpha}}_{=:I_{2}},$ (A.11)
@@
A.3.2 Bounding the first term $I_{1}$
@@
Recall that $\|\mathcal{C}_{\pi}^{1/2}{\mathcal{C}}_{\alpha}^{-1/2}\|_{{\mathcal{L}}({\mathcal{H}})}\leq 1$ and that by Corollary C.9 and Lemma D.2, with confidence $1-\delta$, we have
@@
$\|I_{1}\|_{L^{2}(\pi)}=\lVert I_{\pi}(\widehat{\mathcal{C}}_{\pi}+\alpha\operatorname{Id}_{\mathcal{H}})^{-1}(\widehat{\Upsilon}-\widehat{\mathcal{C}}_{\pi}f_{\alpha})\rVert_{L^{2}(\pi)}$ $\leq 4\;\lVert(\mathcal{C}_{\pi}+\alpha\operatorname{Id}_{\mathcal{H}})^{-1/2}(\widehat{\Upsilon}-\widehat{\mathcal{C}}_{\pi}f_{\alpha})\rVert_{\mathcal{H}}.$ (A.12)
@@
Thus, with (A.12), we have
$\|I_{1}\|_{L^{2}(\pi)}\leq 4\mathopen{}\mathclose{{}\left(\|I_{11}\|_{\mathcal{H}}+\|I_{12}\|_{\mathcal{H}}+\|I_{13}\|_{\mathcal{H}}}\right)$ (A.13)
@@
To bound the norm of $I_{11}$, we apply Bennett’s inequality Proposition C.2 to the independent and identically distributed random variables
@@
In the last step we use (A.3.2). From Proposition C.2, we obtain with confidence $1-\delta$
@@
$\leq c_{11}\mathopen{}\mathclose{{}\left(\frac{\log(2/\delta)}{\sqrt{\alpha}n}+\sqrt{\frac{{\mathcal{N}}(\alpha)\log(2/\delta)}{n}}}\right),$ (A.15)
@@
$\leq c_{12}\sqrt{{\mathcal{N}}(\alpha)}\;\max\mathopen{}\mathclose{{}\left\{\mathopen{}\mathclose{{}\left(\frac{1}{\delta n^{q-1}}}\right)^{1/q}\mathopen{}\mathclose{{}\left(\frac{1}{\alpha{\mathcal{N}}(\alpha)}}\right)^{\frac{q-2}{2q}},\sqrt{\frac{\log(2c_{1}/\delta)}{n}}}\right\},$ (A.16)
@@
$\leq R\max\{1,{\kappa^{2(\nu-1)}}\}\,\alpha^{\min\{\nu,1\}}=c_{13}\,\alpha^{\min\{\nu,1\}},$ (A.17)
where we use Lemma D.1 for the last inequality and set $c_{13}:=R\max\{1,{\kappa^{2(\nu-1)}}\}$.
@@
Collecting the bounds (A.3.2), (A.3.2), (A.3.2), together with (A.13) and taking a union bound gives with confidence $1-\delta$
@@
$\leq c_{\kappa,\nu,q,R}\;\mathopen{}\mathclose{{}\left(\alpha^{\min\{\nu,1\}}+\frac{\log(4/\delta)}{\sqrt{\alpha}n}+\sqrt{\frac{{\mathcal{N}}(\alpha)\log(4/\delta)}{n}}+\sqrt{{\mathcal{N}}(\alpha)}\cdot\eta(\delta,n,\alpha)}\right),$ (A.18)
@@
A.3.3 Bounding the second term $I_{2}$
We write again ${\mathcal{C}}_{\alpha}:=\mathcal{C}_{\pi}+\alpha\operatorname{Id}_{\mathcal{H}}$, $\widehat{{\mathcal{C}}}_{\alpha}:=\widehat{\mathcal{C}}_{\pi}+\alpha\operatorname{Id}_{\mathcal{H}}$ and find with Lemma D.2 and Corollary C.9 with confidence $1-\delta$
@@ A.3.3 Bounding the second term $I_{2}$
$\leq 4R\max\{1,{\kappa^{2(\nu-1)}}\}\,\alpha^{\min\{\nu,1\}}=4c_{13}\,\alpha^{\min\{\nu,1\}},$ (A.19)
where we repeat the computation from (A.3.2).
A.3.4 Final variance bound
Collecting now (A.3.1), (A.3.2) and (A.3.3) and taking a union bound gives the final bound for the variance. With confidence $1-\delta$, we obtain
@@ A.3.4 Final variance bound
A.3.5 Bounding the bias $I_{\pi}f_{\alpha}-f_{\star}$
The bias can be bounded in the same way as in Section A.1.2:
@@ A.3.5 Bounding the bias $I_{\pi}f_{\alpha}-f_{\star}$
A.4 Proof of Corollary 4.4
We first need a standard lemma that allows us to bound the effective dimension based on the eigenvalue decay of $\mathcal{T}_{\pi}$ due to [11, Lemma 11].
Lemma A.1 (Eigenvalue decay).
@@
$\|I_{\pi}\widehat{f}_{\alpha}-f_{\star}\|_{L^{2}(\pi)}$ $\leq c\;\mathopen{}\mathclose{{}\left(\alpha^{\min\{\nu,1\}}+\frac{\log(8/\delta)}{\sqrt{\alpha}n}+\sqrt{\frac{{\mathcal{N}}(\alpha)\log(8/\delta)}{n}}+\sqrt{{\mathcal{N}}(\alpha)}\cdot\eta(\delta,n,\alpha)}\right),$ (A.20)
@@
$\sqrt{{\mathcal{N}}(\alpha(n,\delta))}\cdot\eta(\delta,n,\alpha(n,\delta))\lesssim\sqrt{\frac{\log(8c_{1}/\delta)}{n\alpha(n,\delta)^{p}}}$ (A.21)
for all $n\geq\overline{n}_{0}$, i.e., the subgaussian term asymptotically dominates the regularized Fuk–Nagaev term under the choice $\alpha(n,\delta)$. In fact, note that under $\mathcal{N}(\alpha)\leq\tilde{D}\alpha^{p}$ ensured by Lemma A.1, we have
@@
$\sqrt{{\mathcal{N}}(\alpha)}\cdot\eta(\delta,n,\alpha)\lesssim\max\mathopen{}\mathclose{{}\left\{\mathopen{}\mathclose{{}\left(\frac{1}{\delta n^{q-1}}}\right)^{1/q}\cdot\mathopen{}\mathclose{{}\left(\frac{1}{\alpha}}\right)^{\frac{q-2}{2q}}\mathopen{}\mathclose{{}\left(\frac{\tilde{D}}{\alpha}}\right)^{p/q},\sqrt{\frac{\log(8c_{1}/\delta)}{n\alpha^{p}}}}\right\}.$ (A.22)
We insert the definition of $\alpha=\alpha(n,\delta)$ into (A.22) and isolate the exponents corresponding to $1/n$ in both expressions inside of the above maximum. We obtain the exponent
$\frac{q-1}{q}-\frac{1}{2\min\{\nu,1\}+p}\cdot\mathopen{}\mathclose{{}\left(\frac{q-2}{2q}+\frac{p}{q}}\right)=\frac{2(q-1)(2\min\{\nu,1\}+p)-(q-2+2p)}{2q(2\min\{\nu,1\}+p)}$ (A.23)
@@
$\frac{1}{2}-\frac{p/2}{2\min\{\nu,1\}+p}=\frac{\min\{\nu,1\}}{2\min\{\nu,1\}+p}=\frac{2q\min\{\nu,1\}}{2q(2\min\{\nu,1\}+p)}$ (A.24)
[…]
@@
where we use $\nu\geq 2$, $q\geq 3$ and $p\in(0,1)$, hence proving (A.21).
Consequently, for $n\geq\max\{\widetilde{n}_{0},\overline{n}_{0}\}$ and with Lemma A.1 and (A.21), the bound provided by (A.20) now reduces to
@@
$\lesssim\mathopen{}\mathclose{{}\left(\frac{\log(8c_{1}/\delta)}{n}}\right)^{\frac{\min\{\nu,1\}}{2\min\{\nu,1\}+p}}+\mathopen{}\mathclose{{}\left(\frac{\log(8c_{1}/\delta)}{n}}\right)^{\frac{2\min\{\nu,1\}+p-1/2}{2\min\{\nu,1\}+p}}+\mathopen{}\mathclose{{}\left(\frac{\log(8c_{1}/\delta)}{n}}\right)^{\frac{\min\{\nu,1\}}{2\min\{\nu,1\}+p}}$
@@
Appendix B Excess risk bound for polynomial confidence regime
@@
Corollary B.1 (Polynomial confidence regime).
@@
$\lVert I_{\pi}\widehat{f}_{\alpha_{2}(n,\delta)}-f_{\star}\rVert_{L^{2}(\pi)}$ $\leq c\cdot\alpha_{2}(n,\delta)^{\min\{\nu,1\}}$ (B.1)
@@
$\alpha\geq\mathopen{}\mathclose{{}\left(\frac{1}{n}}\right)^{\frac{1}{2\min\{\nu,1\}-1}},$ (B.2)
@@
$\alpha\geq c_{10}\;\mathopen{}\mathclose{{}\left(\frac{1}{\delta n^{q-1}}}\right)^{\frac{2}{q(2\min\{\nu,1\}+1)}},$ (B.3)
@@
provided conditions $\delta\in D_{2}(n,q)$ (3.1), (B.2), (B.3) and $\alpha\leq\kappa^{2}$ hold. To simplify the conditions for $\alpha$, we observe that for all $q\geq 3$, $\nu\leq 1$ and $\delta\in(0,1)$ we have
@@
As a result, condition (B.3) implies (B.2) with an appropriate constant. To sum up, the regularization parameter needs to satisfy
@@
then (B.1) holds with confidence $1-\delta$.
Appendix C Concentration bounds
@@
C.1 Hoeffding inequality in Hilbert spaces
The following bound is classical and follows as a special case of [63, Theorem 3.5], see also [64, Section A.5.1].
Proposition C.1 (Hoeffding inequality).
@@ C.1 Hoeffding inequality in Hilbert spaces
C.2 Bennett inequality in Hilbert spaces
We now give a version of a Bennett-type inequality going back to [63, Theorem 3.4]. We use the confidence bound version as derived by [4, Lemma 2].
Proposition C.2 (Bennett inequality).

### v3

@@ B.3.1 Bounding the variance $I_{\pi}\widehat{f}_{\alpha}-I_{\pi}f_{\alpha}$
$=\underbrace{I_{\pi}(\widehat{\mathcal{C}}_{\pi}+\alpha\operatorname{Id}_{\mathcal{H}})^{-1}\mathopen{}\mathclose{{\left(\widehat{\Upsilon}-\widehat{\mathcal{C}}_{\pi}f_{\alpha}}}\right)}_{=:I_{1}}+\underbrace{\alpha I_{\pi}(\widehat{\mathcal{C}}_{\pi}+\alpha\operatorname{Id}_{\mathcal{H}})^{-1}f_{\alpha}}_{=:I_{2}},$ (B.11)
@@
B.3.2 Bounding the first term $I_{1}$
@@
Recall that $\|\mathcal{C}_{\pi}^{1/2}{\mathcal{C}}_{\alpha}^{-1/2}\|_{{\mathcal{L}}({\mathcal{H}})}\leq 1$ and that by Corollary D.9 and Lemma E.2, with confidence $1-\delta$, we have
@@
$\|I_{1}\|_{L^{2}(\pi)}=\lVert I_{\pi}(\widehat{\mathcal{C}}_{\pi}+\alpha\operatorname{Id}_{\mathcal{H}})^{-1}(\widehat{\Upsilon}-\widehat{\mathcal{C}}_{\pi}f_{\alpha})\rVert_{L^{2}(\pi)}$ $\leq 4\;\lVert(\mathcal{C}_{\pi}+\alpha\operatorname{Id}_{\mathcal{H}})^{-1/2}(\widehat{\Upsilon}-\widehat{\mathcal{C}}_{\pi}f_{\alpha})\rVert_{\mathcal{H}}.$ (B.12)
@@
Thus, with (B.12), we have
$\|I_{1}\|_{L^{2}(\pi)}\leq 4\mathopen{}\mathclose{{\left(\|I_{11}\|_{\mathcal{H}}+\|I_{12}\|_{\mathcal{H}}+\|I_{13}\|_{\mathcal{H}}}}\right)$ (B.13)
@@
To bound the norm of $I_{11}$, we apply Bennett’s inequality Proposition D.2 to the independent and identically distributed random variables
@@
In the last step we use (B.3.2). From Proposition D.2, we obtain with confidence $1-\delta$
@@
$\leq c_{11}\mathopen{}\mathclose{{\left(\frac{\log(2/\delta)}{\sqrt{\alpha}n}+\sqrt{\frac{{\mathcal{N}}(\alpha)\log(2/\delta)}{n}}}}\right),$ (B.15)
@@
$\leq c_{12}\sqrt{{\mathcal{N}}(\alpha)}\;\max\mathopen{}\mathclose{{\left\{\mathopen{}\mathclose{{\left(\frac{1}{\delta n^{q-1}}}}\right)^{1/q}\mathopen{}\mathclose{{\left(\frac{1}{\alpha{\mathcal{N}}(\alpha)}}}\right)^{\frac{q-2}{2q}},\sqrt{\frac{\log(2c_{1}/\delta)}{n}}}}\right\},$ (B.16)
@@
$\leq R\max\{1,{\kappa^{2(\nu-1)}}\}\,\alpha^{\min\{\nu,1\}}=c_{13}\,\alpha^{\min\{\nu,1\}},$ (B.17)
where we use Lemma E.1 for the last inequality and set $c_{13}:=R\max\{1,{\kappa^{2(\nu-1)}}\}$.
@@
Collecting the bounds (B.3.2), (B.3.2), (B.3.2), together with (B.13) and taking a union bound gives with confidence $1-\delta$
@@
$\leq c_{\kappa,\nu,q,R}\;\mathopen{}\mathclose{{\left(\alpha^{\min\{\nu,1\}}+\frac{\log(4/\delta)}{\sqrt{\alpha}n}+\sqrt{\frac{{\mathcal{N}}(\alpha)\log(4/\delta)}{n}}+\sqrt{{\mathcal{N}}(\alpha)}\cdot\eta(\delta,n,\alpha)}}\right),$ (B.18)
@@
B.3.3 Bounding the second term $I_{2}$
We write again ${\mathcal{C}}_{\alpha}:=\mathcal{C}_{\pi}+\alpha\operatorname{Id}_{\mathcal{H}}$, $\widehat{{\mathcal{C}}}_{\alpha}:=\widehat{\mathcal{C}}_{\pi}+\alpha\operatorname{Id}_{\mathcal{H}}$ and find with Lemma E.2 and Corollary D.9 with confidence $1-\delta$
@@ B.3.3 Bounding the second term $I_{2}$
$\leq 4R\max\{1,{\kappa^{2(\nu-1)}}\}\,\alpha^{\min\{\nu,1\}}=4c_{13}\,\alpha^{\min\{\nu,1\}},$ (B.19)
where we repeat the computation from (B.3.2).
B.3.4 Final variance bound
Collecting now (B.3.1), (B.3.2) and (B.3.3) and taking a union bound gives the final bound for the variance. With confidence $1-\delta$, we obtain
@@ B.3.4 Final variance bound
B.3.5 Bounding the bias $I_{\pi}f_{\alpha}-f_{\star}$
The bias can be bounded in the same way as in Section B.1.2:
@@ B.3.5 Bounding the bias $I_{\pi}f_{\alpha}-f_{\star}$
B.4 Proof of Corollary 4.4
We first need a standard lemma that allows us to bound the effective dimension based on the eigenvalue decay of $\mathcal{T}_{\pi}$ due to [fischer2020sobolev, Lemma 11].
Lemma B.1 (Eigenvalue decay).
@@
$\|I_{\pi}\widehat{f}_{\alpha}-f_{\star}\|_{L^{2}(\pi)}$ $\leq c\;\mathopen{}\mathclose{{\left(\alpha^{\min\{\nu,1\}}+\frac{\log(8/\delta)}{\sqrt{\alpha}n}+\sqrt{\frac{{\mathcal{N}}(\alpha)\log(8/\delta)}{n}}+\sqrt{{\mathcal{N}}(\alpha)}\cdot\eta(\delta,n,\alpha)}}\right),$ (B.20)
@@
$\sqrt{{\mathcal{N}}(\alpha(n,\delta))}\cdot\eta(\delta,n,\alpha(n,\delta))\lesssim\sqrt{\frac{\log(8c_{1}/\delta)}{n\alpha(n,\delta)^{p}}}$ (B.21)
for all $n\geq\overline{n}_{0}$, i.e., the subgaussian term asymptotically dominates the regularized Fuk–Nagaev term under the choice $\alpha(n,\delta)$. In fact, note that under $\mathcal{N}(\alpha)\leq\tilde{D}\alpha^{p}$ ensured by Lemma B.1, we have
@@
$\sqrt{{\mathcal{N}}(\alpha)}\cdot\eta(\delta,n,\alpha)\lesssim\max\mathopen{}\mathclose{{\left\{\mathopen{}\mathclose{{\left(\frac{1}{\delta n^{q-1}}}}\right)^{1/q}\cdot\mathopen{}\mathclose{{\left(\frac{1}{\alpha}}}\right)^{\frac{q-2}{2q}}\mathopen{}\mathclose{{\left(\frac{\tilde{D}}{\alpha}}}\right)^{p/q},\sqrt{\frac{\log(8c_{1}/\delta)}{n\alpha^{p}}}}}\right\}.$ (B.22)
We insert the definition of $\alpha=\alpha(n,\delta)$ into (B.22) and isolate the exponents corresponding to $1/n$ in both expressions inside of the above maximum. We obtain the exponent
$\frac{q-1}{q}-\frac{1}{2\min\{\nu,1\}+p}\cdot\mathopen{}\mathclose{{\left(\frac{q-2}{2q}+\frac{p}{q}}}\right)=\frac{2(q-1)(2\min\{\nu,1\}+p)-(q-2+2p)}{2q(2\min\{\nu,1\}+p)}$ (B.23)
@@
$\frac{1}{2}-\frac{p/2}{2\min\{\nu,1\}+p}=\frac{\min\{\nu,1\}}{2\min\{\nu,1\}+p}=\frac{2q\min\{\nu,1\}}{2q(2\min\{\nu,1\}+p)}$ (B.24)
for the second expression. We show that the difference between the numerator of (B.23) and the numerator of (B.24) is nonnegative. We have
@@
where we use $\nu\geq 2$, $q\geq 3$ and $p\in(0,1)$, hence proving (B.21).
Consequently, for $n\geq\max\{\widetilde{n}_{0},\overline{n}_{0}\}$ and with Lemma B.1 and (B.21), the bound provided by (B.20) now reduces to
@@
$\lesssim\mathopen{}\mathclose{{\left(\frac{\log(8c_{1}/\delta)}{n}}}\right)^{\frac{\min\{\nu,1\}}{2\min\{\nu,1\}+p}}+\mathopen{}\mathclose{{\left(\frac{\log(8c_{1}/\delta)}{n}}}\right)^{\frac{2\min\{\nu,1\}+p-1/2}{2\min\{\nu,1\}+p}}$
$\quad+\mathopen{}\mathclose{{\left(\frac{\log(8c_{1}/\delta)}{n}}}\right)^{\frac{\min\{\nu,1\}}{2\min\{\nu,1\}+p}}$
@@
Appendix C Excess risk bound for polynomial confidence regime
@@
Corollary C.1 (Polynomial confidence regime).
@@
$\lVert I_{\pi}\widehat{f}_{\alpha_{2}(n,\delta)}-f_{\star}\rVert_{L^{2}(\pi)}$ $\leq c\cdot\alpha_{2}(n,\delta)^{\min\{\nu,1\}}$ (C.1)
@@
$\alpha\geq\mathopen{}\mathclose{{\left(\frac{1}{n}}}\right)^{\frac{1}{2\min\{\nu,1\}-1}},$ (C.2)
@@
$\alpha\geq c_{10}\;\mathopen{}\mathclose{{\left(\frac{1}{\delta n^{q-1}}}}\right)^{\frac{2}{q(2\min\{\nu,1\}+1)}},$ (C.3)
@@
provided conditions $\delta\in D_{2}(n,q)$ (3.1), (C.2), (C.3) and $\alpha\leq\kappa^{2}$ hold. To simplify the conditions for $\alpha$, we observe that for all $q\geq 3$, $\nu\leq 1$ and $\delta\in(0,1)$ we have
@@
As a result, condition (C.3) implies (C.2) with an appropriate constant. To sum up, the regularization parameter needs to satisfy
@@
then (C.1) holds with confidence $1-\delta$.
Appendix D Concentration bounds
@@
D.1 Hoeffding inequality in Hilbert spaces
The following bound is classical and follows as a special case of [Pinelis94, Theorem 3.5], see also [Mollenhauer2022Phd, Section A.5.1].
Proposition D.1 (Hoeffding inequality).
@@ D.1 Hoeffding inequality in Hilbert spaces
D.2 Bennett inequality in Hilbert spaces
We now give a version of a Bennett-type inequality going back to [Pinelis94, Theorem 3.4]. We use the confidence bound version as derived by [Smale2007, Lemma 2].
Proposition D.2 (Bennett inequality).